Date: Fri, 17 Oct 1997 09:31:07 -0700 (PDT) From: email@example.com To: firstname.lastname@example.org Cc: email@example.com, firstname.lastname@example.org, email@example.com Subject: 2MASS WG Mtg #133 Minutes IPAC 2MASS Working Group Meeting #133 Minutes 10/14/97 Attendees: R. Beck, R. Cutri, D. Engler, J. Fowler, T. Jarrett, D. Kirkpatrick, G. Kopan, H. McCallon, S. Wheelock, J. White AGENDA 1.) Telescope and Camera Status 2.) Science Team Meeting 3.) 14-Bit Tests 4.) Disk Errors on Rodan 5.) Minor Planet Processing DISCUSSION 1.) Telescope and Camera Status R. Cutri reported the following items: (a.) the southern telescope has arrived safely in Santiago; (b.) laboratory dark and flat sequences have been taken with the southern camera and are on a tape due to arrive soon at IPAC; (c.) the northern telescope's accuracy in pointing and tracking has been good enough since the recent repairs to change the color of many items in the web-based quality reports from red to green; (d.) the increasingly cold weather on Mt. Hopkins has apparently evoked a temperature-dependent astigmatic effect that has altered the relationship between the bands in PSF behavior; (e.) the new calibration strategy has now been in effect at the observatory since last Friday; (f.) the 971005 tape received from the northern observatory employed the new compression standard. 2.) Science Team Meeting R. Cutri announced that the next Science Team Meeting will be in Boston on November 4-5 and will include parameter tuning as one of its main agenda. 3.) 14-Bit Tests After several reruns to get exact duplication of conditions, the 16-bit and 14-bit runs of 970608 have been completed, and some reports have been received concerning differences. All the early results have confirmed, e.g., dark noise increases at the 1% to 3% level, drops in source extraction, more single-band and dubious band combinations (especially J and K with no H), slightly higher coadd noise. The worst hit seems to be in galaxy processing, where T. Jarrett reported magnitude discrepancies in the 2% to 4% range. Analysis is continuing. 4.) Disk Errors on Rodan R. Beck reported that I/O errors have been occurring on the "rawk" disk (where the raw K data are stored for pipeline processing) and have gotten too bad to tolerate; a disk swap will be performed. 5.) Minor Planet Processing J. Fowler reported that the MPCAT subsystem is finally ready for testing. Early runs are encouraging, although an unexpectedly large number of scan have no minor-planet predictions; this is believed to be because most of the scans available online are calibration scans and therefore have areas of only 1/6 square degree, and a large number of scans are dedicated to the same areas, most of which are not very close to the ecliptic. Much more testing will be done, and the format of the associations file will be changed from scan-oriented to observation-night- oriented in order to reduce the number of files with no predictions or associations in them.